An open box is to be made from a square piece of sheet metal by removing a square of side 4 inches from each corner and turning up the edges. If the volume of the box is to be 676 cubic inches, what is the length of the sides of the square of the sheet metal.



Answer by checkley77(12844)

Let x be the sides of the original square piece of sheet metal.
4(x-2*4)^2=676
4(x-8)^2=676
4(x^2-16x+64)=676
4x^2-64x+256-676=0
4x^2-64x-420=0
4(x^2-16x-105)=0
(x-21)(x+5)=0
x-21=0
x=21 ans. for the side measurement of the sheet of metal.
Proof:
4(21-2*4)^2=676
4(21-8)^2=676
4*13^2=676
4*169=676
676=676